A stone after falling freely under the influence of gravity for 1s strikes a glass plate held horizontally. In breaking through the glass plate, the stone losses half the velocity. How far will it travel in next 2s?

Grade:11

1 Answers

Niraj Bishnoi
6 years ago
We know that,
v=u+at
In here,
u/2=g, i.t u=2g
Now, 
s=ut+1/2at2
   =2x2g+5x4
    =60m
So, the stone will move 60m in the next 2s is 60